# Unifying Two Types of Scaling Laws from the Perspective of Conditional Kolmogorov Complexity

### v1

@@ 3.4 Kolmogorov complexity
Kolmogorov complexity has two very important properties:
@@
  • $U(p,y)=x$ means program $p$ runs on universal Turing machine $U$ with input $y$ and outputs $x$.
@@ 3.5 Two Types of Scaling Laws
Within the framework of large language models, Scaling Laws can primarily be categorized into two types:
@@
Kaplan et al. (2020) systematically studied the impact of model parameter scale on language model performance and proposed corresponding Scaling Laws. This type of scaling law primarily focuses on resource optimization during the pre-training phase, with its core being the significant improvement of model performance through increasing key resources such as data volume, model parameter count, and computational power. The proposal of such Scaling Laws not only deepened our understanding of the relationship between model performance and resource investment but also laid an important foundation for the training and optimization of large-scale language models.
Snell et al. (2024) conducted an in-depth study on the feasibility of improving Large Language Model (LLM) performance by increasing computational resources during the inference phase. This research direction was empirically supported by OpenAI’s o1 model released in September 2024. Research shows that significantly increasing computational resources and time investment during the inference phase can effectively improve model performance in complex tasks such as mathematics, programming, and logical reasoning. This finding not only validates the importance of computational scaling during inference but also provides a new technical pathway for optimizing large language model performance. The actual performance of OpenAI’s o1 modelJaech et al. (2024) further confirms that there is a significant positive correlation between inference phase resource investment and model performance, particularly demonstrating important practical value when handling high-complexity tasks.
4 Analysis of Large Language Model Pre-training and Inference from the Perspective of Conditional Kolmogorov Complexity
In this section, we will analyze the pre-training and inference processes of Large Language Models (LLMs) using conditional Kolmogorov complexity, and arrive at the following conclusions:
@@
    The pre-training process of Large Language Models (LLMs) is equivalent to approaching joint Kolmogorov complexity $C(x,y)$ from above.
@@
    Although Kolmogorov complexity is uncomputable, we prove that theoretically there exist total recursive functions or transformer neural networks such that $\lim_{t\to\infty}M^{t}(x,y)=C(x\mid y)$.
@@
    LLMs are constrained by the number of Turing machine execution steps $t$, which prevents them from solving certain problems.
@@
    Theoretically, we have unified the two types of Scaling laws. Both types of Scaling laws work by increasing execution steps $t$ to better approximate conditional Kolmogorov complexity.
4.1 The Relationship Between LLM Pre-training and Conditional Kolmogorov Complexity
@@
$y^{*}=\arg\min_{y}C(x,y)$ (9)
Next, we’ll demonstrate the reasonableness of this formula. According to Zvonkin and Levin (1970), $C(x,y)$ can be further decomposed into the following form:
@@
In fact, the right side of equation 4.1 can be viewed as the Minimum Description Length (MDL) in machine learning. Here, $C(y)$ represents model complexity, while $C(x\mid y)$ represents the encoding length of data under that model. $C(y)$ can be understood as model regularization, while $C(x\mid y)$ reflects how well the model fits the data.
It’s important to note that we measure model complexity using Kolmogorov complexity rather than simply counting the number of model parameters. This approach is quite reasonable. For example, for a transformer network with 10,000 parameters, when all parameters are 1, the network’s complexity is very low; when the parameters are completely random, the network’s complexity increases significantly.
If we only optimize $y$ to minimize $C(x\mid y)$, it may lead to overfitting. For instance, if we directly set $y=x$, then model $y$ becomes exactly equivalent to data $x$, which is essentially just memorization, and in this case $C(y)$ would be very large. A neural network with good generalization ability should start from completely random parameter initialization and gradually learn, during which process both $C(y)$ and $C(x\mid y)$ should decrease gradually.
Below, we will demonstrate through rigorous mathematical derivation that Large Language Model (LLM) pre-training is equivalent to directly approaching joint Kolmogorov complexity $C(x,y)$ from above, and naturally considers model complexity during the training process.
@@
Proof.
Let $p$ be the shortest program describing $y$, and $q$ be the shortest program describing $x$ given $y$. We can construct the following Turing machine: first use $p$ to describe $y$, then mark the end position of $p$ in the Turing machine with the encoding length $l(p)$, then use program $q$ to describe $x$ based on the previously given $y$. The total length of this constructed Turing machine is $C(y)+C(x\mid y)+l(C(y))$. Finally, according to the invariance theorem of Kolmogorov complexity, the above inequality holds. ∎
@@
The pre-training process of LLMs is actually a computable (total recursive) approximation of the right side of equation 4.1.
We use the content from 3.3 to perform lossless compression of $x$. Note that during the lossless compression process, we not only obtain the encoding of $x$ based on $y$ but also obtain the parameters of model $y$.
@@
  • Arithmetic coding $z_{n}$
  • Code information $F$
  • Required number of decoding iterations $d$
$z_{n}$ and $F$ are computable approximations (from above) of $C(y)+C(x\mid y)+O(1)$. $l(d)$ is a computable approximation (from above) of $l(C(y))$. The smaller the encoding length of $z_{n}$, the closer it gets to the upper bound of $C(x,y)$. In other words, the pre-training process of LLMs is essentially searching for a model $y$ of moderate complexity such that, given $y$, the encoding length of input data $x$ is minimized. This objective reflects the core idea of the Minimum Description Length (MDL) principle, which seeks optimal balance between model complexity and data fitting.
4.2 The Relationship Between LLM Inference and Conditional Kolmogorov Complexity
We appropriately rewrote equation 9 and applied it to infinitely many $x$, obtaining the following expression:
@@ 4.2 The Relationship Between LLM Inference and Conditional Kolmogorov Complexity
In the framework of large language models, $x$ here represents a string concatenated from $(k,c)$, where $k$ is our input (prompt) and $c$ is our expected output (response). We conjecture that $C(x\mid y^{*})$ represents the ideal state of artificial general intelligence (AGI), as it can perform lossless compression on any $x$ we care about, indicating it has predictive capability for any $x$. Note that if $x$ cannot be losslessly compressed, it means $x$ is "random" and lacks regularity, therefore such $x$ are not our targets of interest.
@@
There exists a total recursive function $\phi(t,x,y)$ that is monotonically decreasing with $t$, such that $\lim_{t\to\infty}\phi(t,x,y)=C(x\mid y)$.
@@
Corollary 4.4.
Theoretically, there exists a series of transformer neural networks $M^{t}(x,y)$, such that $\lim_{t\to\infty}M^{t}(x,y)=C(x\mid y)$, where $t$ represents the number of Turing machine execution steps.
In Roberts (2024), the authors proved that transformers are Turing complete under infinite precision rational numbers. Therefore, we can construct transformer $M^{t}(x,y)$ to simulate $\phi(t,x,y)$. According to transformer properties, the relationship between $t$ and input length $l(x)$ is of polynomial complexity. Note that we currently cannot prove whether the large language model obtained using the content from 3.3 can infinitely approximate $C(x\mid y)$.
Corollary 4.5.
Given $x$ and $t$, we cannot determine whether there exists $M^{t}(x,y)=C(x\mid y)$.
Corollary 4.6.
For any $\epsilon>0$ and each $t^{{}^{\prime}}$, there exist infinitely many $x$ such that there is always an $\epsilon$ error between $M^{t^{{}^{\prime}}}(x,y)$ and $C(x\mid y)$.
4.4 tells us that theoretically, we can construct a series of transformer networks to gradually approximate $C(x\mid y)$ (i.e., AGI). However, 4.5 and 4.6 indicate that in practice, we cannot precisely determine how close we are to $C(x\mid y)$. Particularly, 4.6 further shows that there will always be some $x$ that have a certain error with $C(x\mid y)$. Unfortunately, we currently cannot determine whether these $x$ with errors are all truly $x$ that we care about (since some $x$ might be completely meaningless random strings). However, we will prove through contradiction that there indeed exist some $x$ that we care about where, due to the time step limitation $t$, there will be significant errors between $C(x\mid y)$ and $M^{t}(x,y)$.
Note that in the following proof, we assume $NP\neq P$. In fact, even without this assumption, as long as we can construct a problem that exceeds polynomial time complexity, the conclusion still holds.
Suppose string $(k,c)=x$ is as follows:
$k$: Given $n$ sets {n sets}, for each set, does there exist a subset whose sum equals exactly 42? $c$: yes,no,no,yes,no......
where {n sets} are $n$ large integer sets, and $c$ is a string of length $n$ composed of "yes" or "no", determined by the content of {n sets}. Obviously, $x$ represents the subset sum problem, which is an NP-complete problem. Since we assume $NP\neq P$, this means there exists no polynomial-time algorithm that can solve this problem.
Therefore, clearly, $C(x\mid y)=C(k\mid y)+l(p)+O(1)$, where $l(p)$ represents the length of program $p$ that solves the subset sum problem. Since no program $p$ of polynomial complexity exists, this means the number of steps $t$ that the program describing $C(x\mid y)$ executes on the Turing machine exceeds polynomial complexity. Therefore, $C(x\mid y)\neq M^{t}(x,y)$. In conclusion, we have found a meaningful string $x$ where there exists a significant error between $C(x\mid y)$ and $M^{t}(x,y)$.
To better approximate $C(x\mid y)$, we need to increase the value of $t$ in $M^{t}(x,y)$. In transformer structures, there are two ways to increase $t$ in $M^{t}(x,y)$:
  • Increase the model’s parameter count: As the model’s parameter count increases, under the same $l(x)$, $t$ in $M^{t}(x,y)$ becomes larger.
  • Increase the encoding process steps: For example, introduce intermediate reasoning path $m$ in $x=(k,c)$, so that the encoding of the first character of $c$ is not generated directly from $k$, but from $(k,m)$.
It is not difficult to observe that the first method corresponds to the pre-training scaling laws while the second method corresponds to the inference scaling laws. The second method has multiple implementations, such as: chain-of-thought ($CoT$)Wei et al. (2022), designing complex agent systems on LLMsAnthropic (2024), and OpenAI’s $o1$Jaech et al. (2024), among others. These can all be considered implementations of the second method. They expand the value of $t$ by increasing the number of output tokens. Whether it is the first or the second method, both essentially approximate $C(x\mid y)$ by increasing $t$.
Finally, we raise a question: Are RNN-type linear structures, such as MambaGu and Dao (2023) and RWKVPeng et al. (2023), truly efficient? Despite the polynomial-time complexity of transformers, it is precisely this polynomial-time complexity that enables them to approximate conditional Kolmogorov complexity more rapidly. In contrast, Mamba and RWKV may be less efficient in approximating conditional Kolmogorov complexity.
5 Conclusion
In this paper, we first demonstrate that the pretraining process of large language models is a computable (total recursive) approximation of the upper bound of joint Kolmogorov complexity. Optimizing joint Kolmogorov complexity inherently carries a regularization effect. Next, we prove that, under the condition of infinite precision rational numbers, there exists a series of transformer neural networks such that $\lim_{t\to\infty}M^{t}(x,y)=C(x\mid y)$. Theoretically, to better approximate $C(x\mid y)$, we need to increase $t$ in $M^{t}(x,y)$. The two types of scaling laws represent different methods of increasing $t$.

### v2

@@ 3.4 Kolmogorov complexity
We emphasize two important properties here for Kolmogorov complexity:
@@
  • $U(p,y)=x$ means program $p$ runs on universal Turing machine $U$ with input $y$ and it outputs $x$.
@@ 3.5 Two Types of Scaling Laws
The Scaling Laws in LLMs can primarily be categorized into two types:
@@
Kaplan et al. (2020) systematically studied the impact of model parameter scale on language model performance and proposed corresponding Scaling Laws. It focuses on resource optimization during the pre-training process, in other words, to improve model performance through increasing key resources such as data volume, model parameter count, and computational power. The proposal of such Scaling Laws also laid an important foundation for training and optimization of LLMs.
Snell et al. (2024) conducted an in-depth study on the feasibility of improving LLMs performance by increasing computational resources during the inference process. This research direction was empirically supported by OpenAI’s o1 model released in September 2024. Research shows that significantly increasing computational resources and time investment during the inference process can effectively improve model performance in complex tasks such as mathematics, programming, and logical reasoning. This finding not only validates the importance of computational scaling during inference but also provides a new technical pathway for optimizing LLMs performance. The outstanding performance of OpenAI’s o1 model Jaech et al. (2024) further confirms that there is a significant positive correlation between inference process resource investment and model performance, particularly demonstrating important practical value when handling high-complexity tasks.
4 Analysis of LLMs Pre-training and Inference from the Perspective of Conditional Kolmogorov Complexity
In this section, we will analyze the pre-training and inference processes of LLMs using conditional Kolmogorov complexity, and reach the following conclusions:
@@ 4 Analysis of LLMs Pre-training and Inference from the Perspective of Conditional Kolmogorov Complexity
    The pre-training process of LLMs is equivalent to approximating joint Kolmogorov complexity $C(x,y)$.
@@
    Although Kolmogorov complexity is uncomputable, we prove that theoretically there exist total recursive functions or decoder-only transformer such that $\lim_{t\to\infty}M(t,x,y)=C(x\mid y)$.
@@
    LLMs are constrained by the number of Turing machine execution steps, which prevents them from solving certain problems.
@@
    Theoretically, we have unified the two types of Scaling Laws. Both types of Scaling Laws work by increasing the number of Turing machine execution steps to better approximate conditional Kolmogorov complexity.
4.1 The Relationship Between LLMs Pre-training and Conditional Kolmogorov Complexity
@@
$\min_{y}C(x,y)$ (9)
Next, we’ll demonstrate the reasonableness of this formula. According to Zvonkin and Levin’s (1970) work, $C(x,y)$ can be further decomposed into the following form:
@@
In fact, the right side of Section 4.1 can be viewed as the MDL in machine learning. Here, $C(y)$ represents model complexity, while $C(x\mid y)$ represents the encoding length of data under that model. $C(y)$ can be viewed as model regularization, while $C(x\mid y)$ reflects how well the model fits the data.
It’s important to note that we measure model complexity using Kolmogorov complexity rather than simply counting the number of model parameters. This approach is quite reasonable. For example, for a neural network with 10,000 parameters, when all parameters are 1, the network’s complexity is very low; when the parameters are completely random, the network’s complexity increases significantly.
If we solely optimize $y$ to minimize $C(x\mid y)$, it may result in overfitting. For instance, if we directly set $y=x$, then model $y$ becomes exactly equivalent to data $x$, which is essentially memorization. In this case, $C(x\mid y)$ would be $O(1)$, while $C(y)$ would be very large. A neural network with good generalization capabilities should start from completely random parameter initialization and gradually learn, during which process both $C(y)$ and $C(x\mid y)$ should decrease gradually.
Below, we will demonstrate through rigorous mathematical derivation that LLMs pre-training is equivalent to directly approximating joint Kolmogorov complexity $C(x,y)$ upper bound, and naturally considers model complexity during the training process.
@@
The proof is in Section B.1.
@@
The pre-training process of LLMs is actually a computable (total recursive) approximation of the right side of Theorem 4.1.
We use the content from Section 3.3 to perform lossless compression of $x$. Note that during the lossless compression process, we not only obtain the encoding of $x$ based on $y$ but also obtain the parameters of model $y$.
@@
  • Arithmetic coding $z_{n}$.
  • Program information $F$.
  • Required number of decoding iterations $d$.
Obviously, $C(y)+C(x\mid y)+O(1)\leq|z_{n}|+|F|+|d|$ and $2l(C(y))+O(1)\leq 2l(|z_{n}|+|F|+|d|)$. The smaller the encoding length of $z_{n}$, the closer it gets to the upper bound of $C(x,y)$. In other words, the pre-training process of LLMs is essentially searching for a model $y$ of moderate complexity such that, given $y$, the encoding length of input data $x$ is minimized. This objective reflects the core idea of the MDL principle, which seeks optimal balance between model complexity and data fitting.
4.2 The Relationship Between LLMs Inference and Conditional Kolmogorov Complexity
We rewrote Equation 9 and applied it to infinite $x$, obtaining the following expression:
@@ 4.2 The Relationship Between LLMs Inference and Conditional Kolmogorov Complexity
In the LLMs, $x$ here represents a string concatenated from $(k,r)$, where $k$ is our input (prompt) and $r$ is our expected output (response). We conjecture that $C(x\mid y^{*})$ represents the ideal state of artificial general intelligence (AGI), as it can perform lossless compression on any $x$ we care about, indicating it has predictive capability for any $x$. Note that if $x$ cannot be losslessly compressed, it means $x$ is "random" and lacks regularity, therefore such $x$ are not our targets of interest.
@@
For each $x$, there exists a total recursive function $\phi(t,x,y)$ that is monotonically decreasing with $t$, such that $\lim_{t\to\infty}\phi(t,x,y)=C(x\mid y)$.
The proof is in Section B.2.
Corollary 4.2.
Theoretically, there exists a series of decoder-only transformer $M(t,x,y)$, such that $\lim_{t\to\infty}M(t,x,y)=C(x\mid y)$.
Roberts (2024) proved that decoder-only transformer are Turing complete under infinite-precision rational numbers and hardmax attention mechanisms. Under this assumption, we can construct decoder-only transformer $M(t,x,y)$ to simulate $\phi(t,x,y)$. Note that we currently cannot prove whether the LLMs and arithmetic coding obtained using the content from Section 3.3 can infinitely approximate $C(x\mid y)$.
Corollary 4.3.
Given $x$ and $t$, we cannot determine whether there exists $M(t,x,y)=C(x\mid y)$.
Corollary 4.4.
For any $\epsilon>0$ and each $t$, there exist infinite $x$ such that there is always an $\epsilon$ error between $M(t,x,y)$ and $C(x\mid y)$.
Corollary 4.2 tells us that theoretically, we can construct a series of decoder-only transformer to gradually approximate $C(x\mid y)$ (i.e., AGI). However, Corollary 4.3 and Corollary 4.4 indicate that in practice, we cannot precisely determine how close we are to $C(x\mid y)$. Particularly, Corollary 4.4 further shows that there will always be some $x$ that have a certain error with $C(x\mid y)$. Unfortunately, we currently cannot determine whether these $x$ with errors are indeed $x$ that we care about (since some $x$ might be completely meaningless random strings). However, When we use practically applied decoder-only transformer $m(x)$ to approximate $C(x\mid y)$, the following interesting corollaries arise.
Corollary 4.5.
For a practically applied decoder-only transformer $m(x)$, there exist meaningful $x$ such that $m(x)\neq C(x\mid y)$.
Note that in the following proof, we assume $NP\neq P$. In fact, even without this assumption, as long as we can construct a problem that can not be solved in polynomial time, the conclusion still holds.
Suppose string $(k,r)=x$ is as follows:
$k$: Given $n$ sets {n sets}, for each set, does there exist a subset whose sum equals exactly 42? $r$: yes,no,no,yes,no......
where {n sets} are $n$ integer sets, and $c$ is a string of length $n$ composed of "yes" or "no", determined by the content of {n sets}. Obviously, $x$ represents the Subset Sum Problem, which is a NP-complete problem. Since we assume $NP\neq P$, this means there exists no polynomial-time Turing machine that can solve this problem.
For a decoder-only transformer, $m(x)$ is a Turing machine with polynomial-time complexity. Assuming for any $x$ (here referring to the use of different sets of integers {n sets}), if $m(x)=C(x\mid y)$, then we have found a Turing machine $m(x)$ with polynomial-time complexity that solves the Subset Sum Problem, leading to a contradiction. Therefore, there exist some $x$ for which $m(x)\neq C(x\mid y)$.
As shown in Theorem 4.2, in order to better approximate the conditional Kolmogorov complexity $C(x\mid y)$, we need to find Turing machines that execute more steps. In decoder-only transformer, there are two ways to increase execution steps, they are illustrated in Figure 1:
  • Increase the model’s parameter count: As the model’s parameter count increases, under the same $l(x)$, execution steps becomes larger.
  • Increase the encoding process steps: For example, introduce intermediate reasoning path $m$ in $x=(k,r)$, so that the encoding of the first character of $r$ is not generated directly from $k$, but from $(k,m)$.
It is easy to observe that the first method corresponds to the pre-training Scaling Laws while the second method corresponds to the inference Scaling Laws. The second method has multiple implementations, such as chain-of-thought Wei et al. (2022), designing complex agent systems based on LLMs Anthropic (2024), adding a "wait" command to extend the model’s thinking time Muennighoff et al. (2025); xjdr-alt (2024), and leveraging advanced models like OpenAI’s o1 Jaech et al. (2024) and DeepSeek’s R1 Guo et al. (2025), among others. These can all be considered implementations of the second method. They expand the value of execution steps by increasing the number of intermediate tokens.
Finally, we raise a question: Are RNN-type linear structures, such as Mamba Gu and Dao (2023) and RWKV Peng et al. (2023), truly efficient? Despite the decoder-only transformer’s polynomial-time complexity, its complexity does accelerate the approximation of conditional Kolmogorov complexity. In contrast, Mamba and RWKV may be less efficient in approximating conditional Kolmogorov complexity.
5 Conclusion
In this paper, we first demonstrate that the pretraining process of LLMs is a computable (total recursive) approximation of the upper bound of joint Kolmogorov complexity. Optimizing joint Kolmogorov complexity inherently carries a regularization effect. Next, we prove that, under the condition of infinite-precision rational numbers and hardmax attention mechanisms, there exists a series of decoder-only transformer such that $\lim_{t\to\infty}M(t,x,y)=C(x\mid y)$. Theoretically, to better approximate $C(x\mid y)$, we need to increase execution steps in $M(t,x,y)$. The two types of Scaling Laws represent different methods of increasing execution steps.
6 Limitations
Our analysis is subject to several notable limitations. First, the existence of $y^{*}$ in Equation 12 remains unproven, which poses a challenge to the theoretical foundation of our approach. Second, the current training methodologies for LLMs may not sufficiently approximate the theoretical construct $M(t,x,y)$. Lastly, the realization of $M(t,x,y)$ depends on infinite-precision rational numbers and hardmax attention mechanisms.
Appendix A Example of Arithmetic Coding
Consider a symbol set $\mathcal{S}=\{a,b,c,d,e\}$, and we need to encode the symbol sequence "bab".
$(0.01001)_{2}=(0.28125)_{10}$
Appendix B Proof of Theorem
B.1 Proof of Theorem 4.1
Theorem 4.1.
$C(x,y)\leq$ $C(y)+C(x\mid y)$
$+2l(C(y))+O(1)$ (13)
Proof.
Let $p$ be the shortest program describing $y$, and $q$ be the shortest program describing $x$ given $y$. We can construct the following Turing machine: first use $p$ to describe $y$, then mark the end position of $p$ in the Turing machine with the encoding length $l(p)$, then use program $q$ to describe $x$ based on the previously given $y$. Therefore, exsit a program $\overline{l(p)}pq$ to describe $\langle x,y\rangle$, where $\overline{l(p)}$ represents a prefix code($\overline{x}=1^{l(x)}0x$). Finally, according to the invariance theorem of Kolmogorov complexity $C(x,y)\leq C(y)+C(x\mid y)+2l(C(y))+O(1)$. ∎
B.2 Proof of Theorem 4.2
Theorem 4.2.
For each $x$, there exists a total recursive function $\phi(t,x,y)$ that is monotonically decreasing with $t$, such that $\lim_{t\to\infty}\phi(t,x,y)=C(x\mid y)$.